### v1

# From Small to Large: A Graph Convolutional Network Approach for Solving Assortment Optimization Problems

Abstract
Assortment optimization involves selecting a subset of substitutable products (subject to certain constraints) to maximize the expected revenue. It is a classic problem in revenue management and finds applications across various industries. However, the problem is usually NP-hard due to its combinatorial and non-linear nature. In this work, we explore how graph concolutional networks (GCNs) can be leveraged to efficiently solve constrained assortment optimization under the mixed multinomial logit choice model. We first develop a graph representation of the assortment problem, then train a GCN to learn the patterns of optimal assortments, and lastly propose two inference policies based on the GCN’s output. Due to the GCN’s inherent ability to generalize across inputs of varying sizes, we can use a GCN trained on small-scale instances to facilitate large-scale instances. Extensive numerical experiments demonstrate that given a GCN trained on small-scale instances (e.g., with 20 products), the proposed policies can achieve superior performance (90%+ optimality) on large-scale instances (with up to 2,000 products) within seconds, which outperform existing heuristic policies in both performance and efficiency. Furthermore, we extend our framework to a model-free setting where the underlying choice model is unknown but transaction data is available. We also conduct numerical experiments to demonstrate the effectiveness and efficiency of our proposed policies in this setting.
Keywords: graph convolutional network; assortment optimization; choice model.
1 Introduction
Assortment optimization is a fundamental problem in revenue management with broad applications across various industries, including airline, hospitality, retailing, and online advertisement. The objective is to select a subset of substitutable products in order to maximize the expected revenue. In practice, there are often constraints on the set of products that can be selected. For example, given a query keyword, an e-commerce platform needs to choose a subset of substitutable products to display on the user interface with limited space. In some other cases, a product can be chosen only if another product is chosen. The assortment optimization problem is usually computationally intractable due to its combinatorial and non-linear nature. For instance, Bront et al. (2009) show that the unconstrained assortment problem under the mixed multinomial logit model is NP-hard even when there are only two customer types, and Désir et al. (2022) demonstrate that the assortment problem under the multinomial logit model is NP-hard when there exists a capacity constraint. However, real-world applications often require frequently solving (personalized) assortment problems in response to customer queries (see, e.g., Golrezaei et al. 2014, Agrawal et al. 2019, and Li et al. 2025). For instance, Amazon processes approximately 850 customer queries per second (Statista 2025), implying that each assortment problem needs to be solved in milliseconds. Therefore, developing fast and reliable algorithms for assortment optimization is of paramount importance for practical deployment.
In recent years, neural networks have shown remarkable potential in various combinatorial optimization tasks such as vehicle routing (Nazari et al. 2018), scheduling (Khalil et al. 2016), and matching (Li et al. 2019). Motivated by these advances, we explore how neural networks can be exploited to facilitate the solution of constrained assortment optimization problems. The high-level idea is to leverage neural networks to learn the patterns of optimal solutions using some training data, and then design some efficient and well-performing policies. In particular, we are interested in answering the following questions:
  1. 1.
    With the knowledge of the choice model, how can we efficiently derive high-quality solutions using neural networks?
  2. 2.
    How does the neural network-based approach perform as the problem size scales up?
  3. 3.
    Without the knowledge of the choice model, how should the policy be adapted to leverage transaction data for assortment optimization?
To address the first question, we consider a model-based case, where both the underlying choice model and its parameters are assumed to be known. We choose the mixed MNL (MMNL) model as the underlying choice model to demonstrate our result. The reason for choosing the MMNL model is because it can approximate a large class of choice models to any degree of accuracy under mild conditions (McFadden and Train 2000). In this case, our objective is to train a neural network to learn the patterns of optimal assortment, and then design inference policies based on the trained neural network. Specifically, we adopt the graph convolutional network (GCN), which has been proven effective for many optimization tasks (see Cappart et al. 2023 for a review). Since the GCN only admits inputs represented by graphs, we provide the graph representation for the constrained assortment problems under the MMNL model. Then, we generate a training dataset by solving a modest number of problem instances to obtain their optimal assortments, which serve as supervision labels. Since the task is to predict whether each product is in the optimal assortment or not, the training process is similar to that of a classification problem. Once trained, given the graph representation of a problem instance to solve, the GCN can output the probability of each product being in the optimal assortment.
Given the probabilities, we still need inference policies to derive solutions. In this work, we first propose an index policy that sorts all products in a descending order by their indices and then evaluates all feasible assortments consisting of the top-$n$ products according to the indices for $n=1,2,\dots,N$, where $N$ is the number of products. Such an approach is inspired by the well-studied “revenue-order policy” in the literature, which can be interpreted as an index policy based on prices. In our work, we implement the index policy based on the probabilities predicted by the GCN, which is referred to as the “GCN” policy. This policy is computationally efficient because the number of evaluated assortments is no greater than $N$, which is significantly smaller than the total number of assortments $12^{N}-$ 1. Moreover, we also consider a “GCN-LS” policy that adopts the local-search technique in Gallego et al. (2024) to further improve the solution derived by the GCN policy.
To address the second question, we exploit a key advantage of the GCN: its inherent ability to generalize across problem instances of varying sizes. Owing to the weight-sharing mechanism across nodes in the graph, a GCN trained on small-scale instances can naturally accommodate large-scale instances. In order to investigate its generalization capability, we conduct extensive numerical experiments. In particular, we use 1,000 samples with 20 products to train the GCN, and then test it on large-scale instances with up to 2,000 products. We find that the GCN policy can achieve excellent performance (more than 90% competitive ratio) in milliseconds, and hence outperform existing heuristic policies in both performance and efficiency. Moreover, compared with the GCN policy, the GCN-LS policy can further improve the performance but requires a longer computation time. Lastly, we also test the performance of the proposed policies on testing samples with different numbers of products, customer types, and constraints, and provide additional insights.
To address the last question, we investigate a model-free case where the underlying choice model is unknown but historical transaction data is available to guide policy learning. In this case, we first use the transaction data to train one GCN, which is referred to as the “choice-GCN”, to approximate the choice probabilities. Then, we generate constrained assortment problem instances, and use the choice-GCN to evaluate all feasible assortments to derive approximated solutions. Based on such solutions, we train another GCN called “solution-GCN”, which is similar to the GCN for the model-based case. Based on the probabilities predicted by the solution-GCN, we use the same inference policies as the model-based case, except that the expected revenue of each candidate assortment is evaluated by the choice-GCN rather than the underlying chioce model. Based on the numerical results, the GCN policy and the GCN-LS policy can achieve 80%+ and 90%+ of the optimal expected revenue, respectively, implying the effectiveness and the efficiency of the proposed policies. Moreover, we highlight that the choice-GCN can be replaced by other estimation models in the literature (e.g., Aouad and Désir 2022, Wang et al. 2023b, and Liu et al. 2025), and then a similar approach can still result in high-quality solutions as long as the estimation model can provide a high-quality choice probability prediction for assortments of varying sizes.
To summarize, to the best of our knowledge, our work presents the first learning-based framework for solving the constrained assortment optimization problem. To address the challenge of scalability, we introduce the GCN which inherently accommodates inputs of varying sizes due to its weight-sharing mechanism. Moreover, the GCN-based approach can address both the known and unknown choice model cases. Our numerical experiments demonstrate that the GCN-based approach has great potential in providing fast and high-quality solutions for constrained assortment optimization problems.
The remainder of this paper is organized as follows. In the rest of this section, we review literature related to our work. In Section 2, we formally introduce the constrained assortment optimization problem. In Section 3, we propose the GCN policy and the GCN-LS policy for the model-based case. In Section 4, we compare our policies with several existing policies and provide additional insights through extensive numerical experiments. In Section 5, we provide a general framework for the model-free case, and then test its performance through numerical experiments. We conclude the paper in Section 6.
1.1 Literature Review
This work is broadly related to three streams of research: choice modeling, assortment optimization, and GCN for optimization. In the following, we review these streams respectively.
Choice modeling. Modeling customers’ choice behaviors is fundamental and crucial in revenue management. In order to depict the substitution between products, researchers have proposed various discrete choice models. We refer readers to Gallego and Topaloglu (2019) and Wang (2021) for comprehensive reviews of choice models. In the following, we only review some related works. One of the most well-studied choice model is the multinomial logit (MNL) model proposed in McFadden (1972). It has gained significant attention from researchers in economics and operations research. For instance, Feldman et al. (2022) conduct field experiments on Alibaba platform and show that the policy based on the MNL model can significantly outperform the existing machine learning policy ignoring the substitution effect. Despite its broad applications, the MNL model has also been criticized due to the unrealistic independence of irrelevant alternatives (IIA) property (Hausman and McFadden 1984). In order to address this challenge, McFadden and Train (2000) propose the mixed MNL (MMNL) model, which assumes that the market consists of customers with different latent types, and the behavior of each type is characterized by an MNL model. It is also shown in McFadden and Train (2000) that the MMNL can approximate a large class of choice models to any degree of accuracy under mild conditions. Other popular choice models include the nested logit (NL) model (McFadden 1980), the Markov chain (MC) model (Blanchet et al. 2016), and the exponomial model (Alptekinoğlu and Semple 2016). In this work, we present how to construct graph representations for these choice models.
In recent years, there have been some works using neural networks to approximate customers’ choice probabilities. For example, in order to mitigate the IIA property of the MNL model, Rosenfeld et al. (2020) adopt a permutation-invariant neural network, which can be seen as an edge-free counterpart of GCN, to extract features related to the provided product set such that the resulting choice probabilities can depend on the product set. Following this idea, Aouad and Désir (2022) incorporate a sample-average approximation scheme such that the resulting approximation is consistent with the random utility theory. Wang et al. (2023b) and Liu et al. (2025) adopt the attention mechanism and a variant of GCN, respectively, such that the neural networks can better capture cross-product interactions. In this work, we provide a general framework to use neural networks to represent various choice models but our focus is not on estimating the choice probabilities but rather on the corresponding assortment optimization problems.
Assortment optimization. Assortment optimization is a classic problem in revenue management, and many researchers are devoted to proposing efficient policies for the problem under different choice models. Among many others, the revenue-order policy has gained significant attention from both researchers and practitioners due to its efficiency. For many unconstrained assortment problems, researchers have provided performance guarantees of the revenue-order policy. For example, Talluri and Van Ryzin (2004) show that the revenue-order policy is optimal for the unconstrained assortment optimization problem under the MNL model. Under the MMNL model and the MNL model with network effects, Rusmevichientong et al. (2014) and Wang and Wang (2017) identify some conditions under which the revenue-order policy is optimal. Furthermore, Berbeglia and Joret (2020), Gallego and Berbeglia (2021) and Gallego and Berbeglia (2021) provide performance guarantees for the policy when the choice model satisfies some mild conditions.
As mentioned, the revenue-order policy is a special index policy, and the performance can be further improved if we design suitable indices. For example, Rusmevichientong et al. (2010) establish the optimality of an index policy for the cardinality-constrained problem under the MNL model, and provide an efficient algorithm to compute the indices. Gallego et al. (2015) show that an index policy based on modified revenues is optimal for the unconstrained problem under the general attraction model. In this work, we show that with a properly-designed machine learning approach, the corresponding index policy can perform quite well (even for large-scale problems). In addition, we adopt a local-search policy similar to Gallego et al. (2024) to further improve the solution obtained from the index policy.
There are two recent works that utilize machine learning approaches to solve model-free assortment problems. First, Chen et al. (2023) propose a variant of GCN to approximate the choice probabilities, and then directly implement a local-search-style algorithm based on the trained network. In contrast, based on the choice-GCN, we generate solution samples and train another GCN to learn the pattern of the optimal solution, resulting in a computationally efficient and well-performing policy. Second, Wang et al. (2023a) train a fully connected neural network to approximate the choice probabilities, and then formulate the assortment problem as a non-linear integer programming based on the closed-form expression of the neural network. However, as the number of products increases, the number of learnable weights of the neural network increases linearly, and the computation time of the non-linear integer programming will become prohibitive. In contrast, the number of weights in our proposed GCNs does not increase in the product number, and the inference policies can output high-quality solutions efficiently.
GCN for optimization. As an emerging topic in computer science, learning to optimize typically refers to training neural networks to imitate the behaviors of a (near-)optimal policy. In this area, the efficiency of GCNs has already been demonstrated in several related tasks. Readers are referred to Cappart et al. (2023) for a comprehensive review. In the following, we review some closely related works. One of the pioneering works is Gasse et al. (2019) which adopts the GCN to guide the branch-and-bound algorithm for mixed integer linear programming. In particular, the authors first construct a bipartite graph to represent the constraints and the objective, and then train a GCN to imitate the behaviors of a strong branching rule, which is time-consuming but usually results in less branching operations. Their numerical experiments show that a GCN trained on small-scaled instances can speed up the computation of large-scale instances. Following this idea, Fan et al. (2023) train a GCN to determine the initial basis in the simplex method for large-scale linear programs. Rather than relying on existing heuristics, they directly learn from training samples with the optimal basis as the label. Similarly, Liu et al. (2024) utilize a GCN to determine the pivoting rule of the simplex method. They first propose two prophet policies based on the knowledge of the optimal solution, and then train a GCN to imitate their behaviors based on only the problem parameters and local information related to the current basis. In this work, we also represent constraints by a bipartite graph, and train a GCN to directly learn from optimal solutions. However, since the objective function of the assortment problem is non-linear and non-monotone, we need to modify the bipartite graph based on the problem’s features, and design new inference policies accordingly.
2 Model
In this work, we study the assortment optimization problem under linear constraints. Consider $N$ products with the fixed price vector $\boldsymbol{r}\in\mathbb{R}_{+}^{N}$. Without loss of generality, we assume that the product costs are zero. Let $[N]$ denote the whole set $\{1,2,\dots,N\}$ and 00 denote the no-purchase option. Given an assortment $S\subseteq[N]$ represented by $\boldsymbol{x}_{S}=\left\{x_{S,j}=\mathbbm{1}_{\{j\in S\}}\right\}_{j=1}^{N}$, the choice probability of product $j$ is denoted by $\phi_{j}(\boldsymbol{x}_{S})$, which can be captured by some choice model. Moreover, we consider $M$ constraints represented by $\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}$, where $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ and $\boldsymbol{b}\in\mathbb{R}^{M\times 1}$. Then, the constrained assortment optimization problem can be formulated as a non-linear integer program as follows:
$\max_{\boldsymbol{x}\in\{0,1\}^{N}}\quad$ $\mathcal{R}(\boldsymbol{x}):=\sum_{j=1}^{N}\phi_{j}(\boldsymbol{x})\cdot r_{i}\qquad\mbox{s.t.}\quad\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}.$ (1)
In this work, we first consider a specific choice model — the MMNL model (in Appendix B, we also consider several other choice models). The MMNL model subsumes the well-studied MNL model and can approximate a large class of choice models to any degree of accuracy under mild conditions (McFadden and Train 2000). Under the MMNL model, customers are assumed to have $K$ latent types which are unobservable to the seller, and the proportion of type-$k$ customers is $\alpha_{k}\in\mathbb{R}_{+}$ with $\sum_{k=1}^{K}\alpha_{k}=1$. The purchase behavior of each type is characterized by an MNL model. Specifically, the random utility of product $j$ for type-$k$ customers is $U_{jk}=u_{jk}-r_{j}+\xi_{jk},$ where $u_{jk}-r_{j}$ is the intrinsic utility of product $j$ for type-$k$ customers, and $\xi_{jk}$’s are i.i.d. standard Gumbel random variables capturing consumer-specific idiosyncracies. As for the no-purchase option, we set $u_{0k}=0$ and $r_{0}=0$, and hence its utility for type-$k$ customers is $U_{0k}=\xi_{0k}$. Then, given an assortment represented by $\boldsymbol{x}\in\{0,1\}^{N}$, the choice probability of product $j$ among type-$k$ customers is
$\psi_{jk}(\boldsymbol{x})=\mathbb{P}\left(U_{jk}\cdot x_{j}\geq U_{\ell k}\cdot x_{\ell},\forall\ell\in[N]\cup\{0\}\right)=\frac{x_{j}\exp(u_{jk}-r_{j})}{1+\sum_{\ell=1}^{N}x_{\ell}\exp(u_{\ell k}-r_{\ell})}=\frac{x_{j}v_{jk}}{1+\sum_{\ell=1}^{N}x_{\ell}v_{\ell k}},$
where $v_{jk}:=\exp(u_{ik}-r_{i})$. The overall choice probability of product $j$ is as follows:
$\phi_{j}(\boldsymbol{x})=\sum_{k=1}^{K}\alpha_{k}\cdot\psi_{jk}(\boldsymbol{x})=\sum_{k=1}^{K}\frac{\alpha_{k}x_{j}v_{jk}}{1+\sum_{\ell=1}^{N}x_{\ell}v_{\ell k}}.$
According to Bront et al. (2009), the assortment problem (1) under the MMNL model is NP-hard, even when there are only two customer types and no constraints (i.e., $K=2$ and $M=0$). However, as mentioned, such problems are frequently encountered in practice and need to be solved within a short time. In this work, our goal is to propose efficient algorithms that can produce high-quality solutions to (1).
3 GCN-Based Policy
In this section, we introduce the graph convolutional network (GCN), and discuss how to use it to solve the constrained assortment optimization problems. We start with the graph representation of the assortment problem.
3.1 Graph Representation
[…]
Figure 1: Graph representation of constrained assortment problem under MMNL choice model.
In Figure 1, the left-hand side represents the MMNL choice model, and the right-hand side represents the constraints $\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}$. Let $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. Then, we specify the feature vectors. Let $\boldsymbol{Y}^{S}$, $\boldsymbol{Y}^{P}$, and $\boldsymbol{Y}^{C}$ denote the feature matrices of customer nodes, product nodes, and constraint nodes, respectively. Each node is characterized by a vector with $d^{V}=4$ entries, carrying the information of prices, utilities, market composition, and constraints: The feature vector of each customer node $k$ is $\boldsymbol{Y}^{S}_{k,:}:=[0,0,\alpha_{k},0]$; the feature vector of each product node $j$ is $\boldsymbol{Y}^{P}_{j,:}:=\left[r_{j},\frac{1}{K}\sum_{k=1}^{K}v_{jk},0,1\right]$; the feature vector of each constraint node $i$ is $\boldsymbol{Y}^{C}_{i,:}:=[0,0,0,b_{i}]$. Let $\boldsymbol{Z}^{S}$ and $\boldsymbol{Z}^{C}$ denote the feature matrices of edges related to customer nodes and constraint nodes, respectively. Each edge is characterized by a vector with $d^{E}=2$ entries, carrying information of utilities and constraints: The feature vector of the edge between customer segment $k$ and product $j$ is $\boldsymbol{Z}^{S}_{j,k,:}:=[v_{jk},0]$; the feature vector of the edge between product $j$ and constraint $i$ is $\boldsymbol{Z}^{C}_{i,j,:}:=[0,A_{ij}]$. Moreover, if we consider a contextual model, we can incorporate the contextual information at customer and product nodes, and assign the context weights at the edges between customers and products.
3.2 GCN
With the above graph representation, we can introduce a GCN with two message-passing layers, one dropout layer, and one Sigmoid layer. To facilitate understanding, we illustrate the architecture of the GCN in Figure 2. In the following, we explain the computation details within each layer.
Figure 2: Illustration of GCN network.
First message-passing layer. The first message-passing layer in Figure 2 maps the graph in Figure 1 into a new graph with updated node features and the same edge features. The detailed computation process for the first message-passing layer is as follows:
  1. 1.
    Message construction and local information mapping: In this step, the message-passing layer extracts the information of each node to be passed to its neighbors. Inspired by the bi-directional message-passing layer in Gasse et al. (2019), we also use different extraction matrices for different directions. Specifically, for the customer-to-product direction, the message passed from customer node $k$ to product node $j$ and the mapped local information of product node $j$ are as follows:
    $\boldsymbol{m}_{kj}^{1}:=\text{ReLU}(\boldsymbol{Y}^{S}_{k,:}\boldsymbol{W}_{n}^{1}+\boldsymbol{Z}^{S}_{j,k,:}\boldsymbol{W}_{e}^{1})+\epsilon\boldsymbol{1}\in\mathbb{R}^{1\times d_{\text{mid}}},\qquad\boldsymbol{l}_{j}^{1}:=\boldsymbol{Y}^{P}_{j,:}\boldsymbol{W}_{n}^{1}\in\mathbb{R}^{1\times d_{\text{mid}}},$
    where $\text{ReLU}(\boldsymbol{x})=\max\{\boldsymbol{x},\boldsymbol{0}\}$ is an element-wise maximization between $\boldsymbol{x}$ and $\boldsymbol{0}$, $d_{\text{mid}}$ is a hyperparameter which we set to 32 in our paper, and $\boldsymbol{W}_{n}^{1}\in\mathbb{R}^{d^{V}\times d_{\text{mid}}}$ and $\boldsymbol{W}_{e}^{1}\in\mathbb{R}^{d^{E}\times d_{\text{mid}}}$ are (learnable) feature extraction matrices. The value of $\epsilon$ in the last term is chosen to be a small positive constant (e.g., $10^{-7}$) to make sure that the message is always positive. Similarly, the messages passed from product node $j$ to customer node $k$, from product node $j$ to constraint node $i$, from constraint node $i$ to product node $j$, and the corresponding mapped local information are as follows:
    $\boldsymbol{m}_{jk}^{2}$ $:=\text{ReLU}(\boldsymbol{Y}^{P}_{j,:}\boldsymbol{W}_{n}^{2}+\boldsymbol{Z}^{S}_{j,k,:}\boldsymbol{W}_{e}^{2})+\epsilon\boldsymbol{1}\in\mathbb{R}^{1\times d_{\text{mid}}},\qquad\boldsymbol{l}_{k}^{2}:=\boldsymbol{Y}^{S}_{k,:}\boldsymbol{W}^{2}_{n}\in\mathbb{R}^{1\times d_{\text{mid}}}$
    $\boldsymbol{m}_{ji}^{3}$ $:=\text{ReLU}(\boldsymbol{Y}^{P}_{j,:}\boldsymbol{W}_{n}^{3}+\boldsymbol{Z}^{C}_{i,j,:}\boldsymbol{W}_{e}^{3})+\epsilon\boldsymbol{1}\in\mathbb{R}^{1\times d_{\text{mid}}},\qquad\boldsymbol{l}^{3}_{i}:=\boldsymbol{Y}^{C}_{i,:}\boldsymbol{W}_{n}^{3}\in\mathbb{R}^{1\times d_{\text{mid}}}$
    $\boldsymbol{m}_{ij}^{4}$ $:=\text{ReLU}(\boldsymbol{Y}^{C}_{i,:}\boldsymbol{W}_{n}^{4}+\boldsymbol{Z}^{C}_{i,j,:}\boldsymbol{W}_{e}^{4})+\epsilon\boldsymbol{1}\in\mathbb{R}^{1\times d_{\text{mid}}},\qquad\boldsymbol{l}_{j}^{4}:=\boldsymbol{Y}^{P}_{j,:}\boldsymbol{W}^{4}_{n}\in\mathbb{R}^{1\times d_{\text{mid}}},$
    where $\boldsymbol{W}_{n}^{\ell}\in\mathbb{R}^{d^{V}\times d_{\text{mid}}}$ and $\boldsymbol{W}_{e}^{\ell}\in\mathbb{R}^{d^{E}\times d_{\text{mid}}}$ for $\ell=2,3,4$.
  2. 2.
    Message aggregation: For each node, we aggregate its neighbor messages using an element-wise softmax-weighted summation and then combine the aggregated message with its local information. We first define the element-wise softmax aggregation function as follows:
    ($\mathcal{F}\left(\{\boldsymbol{m}_{1},\boldsymbol{m}_{2},\dots,\boldsymbol{m}_{S}\}\right):=\boldsymbol{m}^{\prime}\in\mathbb{R}^{1\times d_{\text{mid}}},\text{ where $m^{\prime}_{\ell}=\sum_{s=1}^{S}\frac{e^{m_{s,\ell}}}{\sum_{s^{\prime}=1}^{S}e^{m_{s^{\prime},\ell}}}\times m_{s,\ell}$}.$
    Then, the aggregated information vectors of customer node $k$, product node $j$, and constraint node $i$ are as follows:
    $\boldsymbol{h}^{S}_{k}$ :=($:=\mathcal{F}\left(\left\{\boldsymbol{m}^{2}_{jk}\right\}_{j=1}^{N}\right)+\boldsymbol{l}_{k}^{2}\in\mathbb{R}^{1\times d_{\text{mid}}}$
    $\boldsymbol{h}^{P}_{j}$ $:=\frac{1}{2}\mathcal{F}\left(\left\{\boldsymbol{m}^{1}_{kj}\right\}_{k=1}^{K}\right)+\frac{1}{2}\boldsymbol{l}^{1}_{j}+\frac{1}{2}\mathcal{F}\left(\left\{\boldsymbol{m}^{4}_{ij}\right\}_{i=1}^{M}\right)+\frac{1}{2}\boldsymbol{l}_{j}^{4}\in\mathbb{R}^{1\times d_{\text{mid}}}$
    $\boldsymbol{h}^{C}_{i}$ :=($:=\mathcal{F}\left(\left\{\boldsymbol{m}^{3}_{ji}\right\}_{j=1}^{N}\right)+\boldsymbol{l}^{3}_{i}\in\mathbb{R}^{1\times d_{\text{mid}}}.$
  3. 3.
    Node update: Lastly, the message-passing layer uses a two-layer fully connected network with an expansion factor of 2 to extract information from the above combined information. In particular, the updated feature matrices are as follows:
    $\boldsymbol{Y}^{S}_{(1),k,:}=\sigma\left(\boldsymbol{h}^{S}_{k}\boldsymbol{T}_{1}+\boldsymbol{b}_{1}\right)\boldsymbol{T}_{2}+\boldsymbol{b}_{2}\in\mathbb{R}^{1\times d_{\text{mid}}}$
    $\boldsymbol{Y}^{P}_{(1),j,:}=\sigma\left(\boldsymbol{h}^{P}_{j}\boldsymbol{T}_{1}+\boldsymbol{b}_{1}\right)\boldsymbol{T}_{2}+\boldsymbol{b}_{2}\in\mathbb{R}^{1\times d_{\text{mid}}}$
    $\boldsymbol{Y}^{C}_{(1),i,:}=\sigma\left(\boldsymbol{h}^{C}_{i}\boldsymbol{T}_{1}+\boldsymbol{b}_{1}\right)\boldsymbol{T}_{2}+\boldsymbol{b}_{2}\in\mathbb{R}^{1\times d_{\text{mid}}}.$
    where $\sigma(\boldsymbol{x}):=1/(1+e^{-\boldsymbol{x}})$ is an element-wise Sigmoid function, $\boldsymbol{T}_{1}\in\mathbb{R}^{d_{\text{mid}}\times 2d_{\text{mid}}}$, $\boldsymbol{b}_{1}\in\mathbb{R}^{1\times 2d_{\text{mid}}}$, $\boldsymbol{T}_{2}\in\mathbb{R}^{2d_{\text{mid}}\times d_{\text{mid}}}$, and $\boldsymbol{b}_{2}\in\mathbb{R}^{1\times d_{\text{mid}}}$.
To summarize, the message-passing layer updates the feature vector of each node based on its local information and its neighbor messages. For each message-passing direction, all nodes share the weight matrix $\boldsymbol{W}_{n}^{\ell}$ and all edges share the weight matrix $\boldsymbol{W}_{e}^{\ell}$. The weights matrices of the third step (i.e., $\boldsymbol{T}_{\ell}$’s, and $\boldsymbol{b}_{\ell}$’s) are shared by all nodes. Such a mechanism is referred to as the “weight-sharing mechanism”, under which the number of learnable weights is independent of the number of nodes.
Dropout layer. The dropout layer in Figure 2 is introduced to mitigate overfitting (Srivastava et al. 2014). Given the dropout probability $p_{D}=0.2$, during the training period, the value of each entry in $\boldsymbol{Y}^{S}_{(1)}$, $\boldsymbol{Y}^{P}_{(1)}$, and $\boldsymbol{Y}^{C}_{(1)}$ will be set to 0 with probability $p_{D}$. The shapes of these matrices will not be affected. During the testing period, we shut down this layer.
Second message-passing layer. Similar to the computation of the first message-passing layer, this layer updates the feature vector of each node based on both the local information and the neighbor information. The only difference is that the input sizes and the output size are $(d_{\text{mid}},d^{E})$ and $1$, respectively. Thus, the resulting matrix $\boldsymbol{Y}^{P}_{(2)}$ is an $N\times 1\times 1$ vector, which can be transformed into an $N\times 1$ vector $\boldsymbol{y}^{P}_{(2)}$.
Sigmoid layer. The Sigmoid layer transforms the predicted value $y^{P}_{(2),j}$ of each product $j$ into the probability $p_{j}:=\sigma\left(y^{P}_{(2),j}\right)$.
Due to the weight-sharing mechanism and the property of the dropout layer and the Sigmoid layer, the number of learnable weights is independent of the number of nodes, such that a GCN trained on small-scale instances can work for large-scale instances. Moreover, the output of the GCN remains the same as we permute the order of nodes, which is a natural requirement of the mapping from the product set to the optimal assortment. Note that if we directly use a fully connected neural network without the weight-sharing mechanism, then the prediction of the network changes as we permute the input order, resulting in unstable outputs.
In order to train the GCN, we collect historical problem instances to generate training data. Each historical problem instance $\ell$ is characterized by $(\boldsymbol{\alpha}_{\ell},\boldsymbol{r}_{\ell},\boldsymbol{U}_{\ell},\boldsymbol{V}_{\ell},\boldsymbol{A}_{\ell},\boldsymbol{b}_{\ell})$, and we use the conic programming approach proposed in Şen et al. (2018) to obtain the optimal solution $\boldsymbol{x}_{\ell}^{*}$. Then, we split the data samples into 80% training set and 20% validation set. In each training epoch, we plug the graph representations of training samples into the GCN to get the predicted probabilities $\boldsymbol{p}$, and then compute the cross-entropy loss by comparing $\boldsymbol{p}$ and $\boldsymbol{x}^{*}$. Similarly, we can compute the cross-entropy loss on the validation set. Then, we use the Adam algorithm to optimize the weights (Kingma and Ba 2014). As for the stopping rule, we continue the training process until the validation loss shows no improvement for a predefined number of epochs.
3.3 Inference Policy
Given the output of the GCN, we still need to design an inference policy to derive a solution. A natural way is to select the largest feasible set consisting of products with largest $p_{i}$’s. However, due to the non-monotone property of the objective, the resulting policy may perform poorly. To address this challenge, we propose an index policy (see Algorithm 1), which sorts all products in a descending order by their indices and then evaluates all feasible assortments consisting of the top-$n$ products for $n=1,2,\dots,N$ according to the indices.
This policy is motivated by the well-studied revenue-order policy (see, e.g., Talluri and Van Ryzin 2004), which can be interpreted as an index policy based on the prices. Moreover, there are other index policies in the literature: For the cardinality-constrained problem under the MNL model, Rusmevichientong et al. (2010) show that the optimal policy is an index policy based on the values $\left(v_{i}\cdot(r_{i}-\lambda)\right)$’s, where $\lambda$ is a hyperparameter; for the unconstrained problem under the GAM model, Gallego et al. (2015) prove the optimality of an index policy based on the modified prices $\left(p_{i}v_{i}/(v_{i}-w_{i})\right)$’s, where $w_{i}$ is the shadow attraction value of product $i$. In this work, we show that the probabilities predicted by the GCN can result in a well-performing index policy, which is referred to as the “GCN” policy. Furthermore, motivated by the local-search policy in Gallego et al. (2024), we also propose another policy that improves the GCN policy by the local-search technique, which is referred to as the “GCN-LS” policy (see Algorithm 3).
4 Numerical Experiments
In this section, we present the numerical experiments. We perform all numerical experiments on a machine with a 2.0 GHz Quad-Core Intel Core i5 CPU and a NVIDIA GeForce RTX 3080Ti GPU, using Gurobi 12.0, Python 3.10, PyTorch 2.7, and CUDA 11.8. The code is run under Ubuntu 22.04. In the following, we compare our proposed policies with three baseline policies as follows (the detailed algorithms are deferred to Appendix A):
  1. 1.
    Revenue-order (RO) policy: We implement the index policy based on the prices $r_{j}$’s.
  2. 2.
    Local-search (LS) policy: We implement the local search policy in Gallego et al. (2024) (see Algorithm 2).
  3. 3.
    CONIC: We implement the conic programming approach in Şen et al. (2018), which is significantly faster than the traditional MILP formulation in Méndez-Díaz et al. (2014). We set the MIPGap threshold as $0.1\%$ and the time limit of the conic program as 600 seconds.
Solution sample generation. In the numerical experiments, given the number of products $N_{\text{train}}$, the number of constraints $M_{\text{train}}$ and the number of customer types $K_{\text{train}}$, we randomly generate each problem instance as follows:
  • To generate the proportions of different customer types, we first sample $K_{\text{train}}$ values $z_{k}$’s from the uniform distribution $\mathcal{U}[0,1]$, and then choose $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{k})$ where $\alpha_{k}=z_{k}/\left(\sum_{k^{\prime}=1}^{K}z_{k^{\prime}}\right)$.
  • Other parameters are generated as follows: $r_{j}\sim\mathcal{U}[3,4]$, $u_{jk}\sim\mathcal{U}[0,1]$, $A_{ij}\sim\mathcal{U}[0,1]$, and $b_{i}\sim\mathcal{U}[5,10]$, where $\mathcal{U}[a,b]$ represents a uniform distribution over the interval $[a,b]$. Then, we can compute $v_{jk}=\exp(u_{jk}-r_{j})$ for any $j$ and $k$.
  • As mentioned, we use the CONIC policy to compute the optimal assortment $\boldsymbol{x}^{*}$, which serves as the label.
Subsequently, each solution sample is of the form $(\boldsymbol{\alpha},\boldsymbol{r},\boldsymbol{U},\boldsymbol{A},\boldsymbol{b},\boldsymbol{V},\boldsymbol{x}^{*})$. We generate $1,000$ solution samples with $K_{\text{train}}=10$ customer types, $N_{\text{train}}=20$ products, and $M_{\text{train}}=5$ constraints, and the total generation time is 558 seconds. We randomly split the solution samples into training set (80%) and validation set (20%), ensuring no overlap between them. Then, we use these solution samples to train the GCN model. We adopt the Adam optimizer in Kingma and Ba (2014), with learning rate 0.001 and batch size 200. We adopt an early stopping rule based on the cross-entropy loss on the validation set. Specifically, if no improvement in the validation loss is observed for 50 epochs, we terminate the training process and retain the model with the lowest validation loss. The training process terminates after 202 epochs, and the total training time is 39.7 seconds. Under the selected model, the cross-entropy loss on the training set is 0.394 and that on the validation set is 0.379.
Testing. In order to test the performance of different policies, we generate $200$ testing instances with $K_{\text{test}}=10$ customer types, $N_{\text{test}}=200\mbox{ or }2,000$ products, and $M_{\text{train}}=5$ constraints. We use the objective computed by the CONIC policy as the baseline. The performance of each policy is measured by the competitive ratio, i.e., the ratio of the expected revenue of the policy to that of the CONIC policy. The numerical results are provided in Table 1.
Table 1: Performance of GCN network and baseline policies when $K_{\text{train}}=K_{\text{test}}=10$, $M_{\text{train}}=M_{\text{test}}=5$, $N_{\text{train}}=20$ and $N_{\text{test}}=200\mbox{ or }2,000$.
Policies $N_{\text{test}}=200$ $N_{\text{test}}=2,000$
Avg. Ratio (std) Avg. Time (s) Avg. Ratio (std) Avg. Time (s)
GCN 91.6% (3.4%) 0.002 89.6% (3.1%) 0.008
GCN-LS 96.8% (1.2%) 0.933 94.6% (1.3%) 61.692
RO 50.6% (6.3%) 0.001 42.7% (5.1%) 0.004
LS 82.7% (4.5%) 0.345 77.0% (3.2%) 15.226
CONIC 100.0% (0.0%) 215.439 100.0% (0.0%) 692.142
According to Table 1, although the GCN is trained on small-scale instances with 20 products, the performance of the GCN policy on instances with 200 or 2,000 products is significantly better than other heuristic policies (i.e., RO and LS). Since the matrix computation of GCN is fast and can be boosted by GPU, the average running time of the GCN policy is in milliseconds, which is negligible. Moreover, combined with the local-search technique, the performance of the GCN policy can be further improved, with a slightly longer computational time.
101010102020202030303030404040405050505060606060%$80$%$100$%$N_{\text{train}}$ompetitive Ratio
10010010010020020020020030030030030040040040040050050050050060606060%$80$%$100$%$N_{\text{test}}$ompetitive Ratio
555510101010151515152020202060606060%$80$%$100$%$K_{\text{test}}$ompetitive Ratio
222244446666888860606060%$80$%$100$%$M_{\text{test}}$ompetitive Ratio
Figure 3: Average competitive ratios of different policies as functions of instance configurations when $K_{\text{train}}=K_{\text{test}}=10$, $M_{\text{train}}=M_{\text{test}}=5$, $N_{\text{train}}=20$, and $N_{\text{test}}=200$. Each curve corresponds to one policy: GCN (red circles), GCN-LS (blue stars), LS (magenta triangles), and RO (cyan squares).
In the remainder of this section, we investigate the impact of instance configurations, including $N_{\text{train}}$, $N_{\text{test}}$, $K_{\text{test}}$ and $M_{\text{test}}$, on the performance of different policies. We fix $K_{\text{train}}=K_{\text{test}}=10$, $M_{\text{train}}=M_{\text{test}}=5$, $N_{\text{train}}=20$, and $N_{\text{test}}=200$ as the base case. Then we change one parameter each time to study the corresponding effect. The results are shown in Figure 3. According to Figure LABEL:fig:N_train, as the product number of training instances $N_{\text{train}}$ increases, the performance of the proposed policies first increases rapidly and then continues to improve gradually. This reflects the increasing generalizability of the GCN as the training instance size grows. According to Figures LABEL:fig:N_test, LABEL:fig:K_test and LABEL:fig:M_test, as $N_{\text{test}}$, $K_{\text{test}}$, or $M_{\text{test}}$ increases, the proposed policies are relatively stable and consistently outperform baseline policies, implying the robustness of the proposed policies.
5 Model-Free Case
In this section, we consider the case where the underlying choice model is unknown and each product is characterized by some contextual information such as quality and price. In this case, we assume that the parameters in the underlying choice model do not change. Given a product assortment $S$ and the corresponding contextual information matrix $\boldsymbol{I}\in\mathbb{R}^{|S|\times F}$ and price vector $\boldsymbol{r}$, the choice probability of product $i\in S$ is denoted by $\Phi(S,\boldsymbol{I},\boldsymbol{r},i)$, which is unknown to the seller. In practice, we have enormous historical transactions, each of which consists of $(S,\boldsymbol{I},\boldsymbol{r})$, and the customer’s choice $\boldsymbol{X}$ is drawn from a multinomial distribution over the items in $S$ with probabilities $\Phi(S,\boldsymbol{I},\boldsymbol{r},j)$ for $j=1,2,\dots,|S|$.
In this case, we propose a three-step framework with two GCNs. We first train a GCN, which is referred to as “choice-GCN”, to approximate the choice probabilities. The choice-GCN can avoid the loss due to the choice model misspecification. Then, we generate many small-scale constrained assortment optimization instances and compute their optimal assortments by evaluating the revenue of all feasible assortments. By doing so, we can generate enough small-scale training samples to train a GCN, which is referred to as “assortment-GCN”. Given the predicted assortment, we can further conduct a local search to improve the performance. The graph representations used for the choice-GCN and the solution-GCN are illustrated in Figure 4.
(a) Choice-GCN
(b) Solution-GCN
Figure 4: Graph representations for choice-GCN and solution-GCN.
Moreover, we would like to highlight that the choice-GCN can be replaced by other neural networks studied in the literature (e.g., Aouad and Désir 2022, Wang et al. 2023b, and Liu et al. 2025), as long as the neural network can provide high-quality choice probability approximation for assortments of varying sizes. Then, following the same three-step framework, the resulting policy can also work well.
In the remainder of this section, we conduct numerical experiments to test the performance of the model-free case. We consider the mixed MNL model as the unknown underlying model and set the number of segments as $K=10$. We set the dimension of the context as $F=5$. In this case, the choice probability can be expressed as $\Phi(S,\boldsymbol{I},\boldsymbol{r},i)=\sum_{k=1}^{K}\alpha_{k}\frac{\exp\left(\sum_{f=1}^{F}c_{k,f}I_{i,f}-r_{i}\right)}{1+\sum_{j\in S}\exp\left(\sum_{f=1}^{F}c_{k,f}I_{j,f}-r_{j}\right)}$, where $\boldsymbol{\alpha}$ is the market proportion vector and $\boldsymbol{C}$ is the weight matrix. Similar to Section 4, we randomly generate $\boldsymbol{\alpha}$ by sampling from uniform variables. Then, we use the same method to generate the context weight vector $\boldsymbol{C}_{k,:}$ of type-$k$ customers. Throughout the numerical experiments, we fix both $\boldsymbol{\alpha}$ and $\boldsymbol{C}$.
Transaction data generation. In order to simulate practical scenarios, we first generate 5,000 transaction samples with $N_{\text{trans}}=10$ products. Similar to Section 4, for each sample, we randomly generate the parameters as follows: $r_{j}\sim\mathcal{U}[3,4]$, $I_{i,f}\sim\mathcal{U}[0,1]$. For each instance, we compute the purchase probabilities $q_{i}$’s computed based on the mixed MNL model, and then generate the choice vector $\boldsymbol{X}\sim\mbox{Multinomial}(|S|;q_{1},q_{2},\dots,q_{|S|})$.
Choice-GCN training. Similarly, we split the data into 80% training data and 20% validation data, and terminate the training algorithm if the validation loss does not improve for 50 epochs. The training process terminates after 134 epochs, and the total training time is 466.5 seconds. Under the selected model, the cross-entropy loss on the training set is 1.412 and that on the validation set is 1.432.
Solution sample generation. Then, we generate 500 small-scale solution samples with 15 products and 5 constraints. The parameters are generated as follows: $r_{j}\sim\mathcal{U}[3,4]$, $I_{i,f}\sim\mathcal{U}[0,1]$, $A_{ij}\sim\mathcal{U}[0,1]$, $b_{i}\sim\mathcal{U}[3,5]$. Based on the trained choice-GCN, we can approximately solve small-scale problem instances by enumerating all feasible assortments. Note that when calling the choice-GCN, we use the batch operation to speed up the computation. The total generation time is 114 minutes.
Solution-GCN training. Similar to the base model, given the solution samples, we can train the assortment-GCN and use it to infer the optimal assortment. Note that since the underlying model is unknown, we instead use the choice-GCN to evaluate the expected revenue of each candidate assortment. When calling the choice-GCN, we also use the batch operation to speed up all policies. In order to compute the baseline revenue of each sample, we run the CONIC policy based on the knowledge of the choice model parameters.
Table 2: Model-free performance of GCN policy and baseline policies when $S_{\text{trans}}=5,000$, $N_{\text{train}}=15$ and $N_{\text{test}}=200$.
Policy Avg. Ratio (std) Avg. Time (s)
GCN 82.7% (4.5%) 0.074
GCN-LS 90.3% (2.6%) 5.233
RO 50.4% (6.2%) 0.056
LS 79.0% (4.9%) 3.568
According to Table 2, for the model-free case, the GCN policy and the GCN-LS policy can achieve 83% and 90% of the optimal expected revenue. Therefore, even without the knowledge of the choice model, the GCN-based approaches can still output high-quality solutions to the assortment optimization problems, outperforming other heuristic approaches.
6 Concluding Remarks
In this work, we explore how to leverage graph convolutional networks to solve constrained assortment problems under the MMNL model. We start with a model-based case, where both the choice model and its parameters are known. In this case, we use a GCN to learn the patterns of optimal solutions. Based on the output of the GCN, we propose two inference policies. Then, we consider the model-free case, where the underlying choice model is unknown. Extensive numerical experiments are conducted to illustrate the effectiveness and efficiency of the proposed policies.
Appendix
Appendix A Algorithms
In this section, we provide algorithm details omitted in the main text.
Algorithm 1 Index Policy
Input: Indices $\tau_{i}$ for $i=1,2,\dots,N$.
Initialize: $\boldsymbol{x}\leftarrow\boldsymbol{0}$, $\boldsymbol{x}^{*}\leftarrow\boldsymbol{0}$, $\mathcal{R}^{*}\leftarrow 0$.
Sort $\tau_{i}$’s in a descending order $\delta$.
for $i=1,2,\dots,N$ do
     Set $x_{j}\leftarrow\mathbbm{1}\{\tau_{j}\geq\tau_{\delta(i)}\}$ for $j=1,2,\dots,N$;
     if $\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}$ is violated then
         Break;
     end if
     Compute the expected revenue $\mathcal{R}\leftarrow\mathcal{R}(\boldsymbol{x})$;
     if $\mathcal{R}>\mathcal{R}^{*}$ then
         Set $\mathcal{R}^{*}\leftarrow\mathcal{R}$ and $\boldsymbol{x}^{*}\leftarrow\boldsymbol{x}$;
     end if
end for
Return: $\boldsymbol{x}^{*}$ and $\mathcal{R}^{*}$.
Algorithm 2 Local-Search Policy
Input: Initial solution $\boldsymbol{x}$. By default, we have $\boldsymbol{x}=\boldsymbol{0}$.
Initialize: $\mathcal{T}(i)\leftarrow 0$ for each $i\in[N]$.
while $\mathcal{T}(i)<1$ for some $i\in[N]$ do
     Set available product set $\mathcal{A}\leftarrow\{j\in[N]\mid\mathcal{T}(i)<1\}$ and the assortment $S=\{j\in[N]\mid x_{j}=1\}$.
     Set $\mbox{ind}_{A}\leftarrow\operatorname*{arg\,max}_{i\in\mathcal{A}\cap\bar{S}}\mathcal{R}(\boldsymbol{x}+\boldsymbol{e}_{i})$ and $\boldsymbol{x}_{A}\leftarrow\boldsymbol{x}+\boldsymbol{e}_{\mbox{ind}_{A}}$.
     if $\mathcal{R}(\boldsymbol{x}_{A})\geq 1.001\times\mathcal{R}(\boldsymbol{x})$ then
         Set $\boldsymbol{x}\leftarrow\boldsymbol{x}_{A}$.
     else
         Set $\mbox{ind}_{D}\leftarrow\operatorname*{arg\,max}_{i\in S}\mathcal{R}(\boldsymbol{x}-\boldsymbol{e}_{i})$ and $\boldsymbol{x}_{D}\leftarrow\boldsymbol{x}-\boldsymbol{e}_{\mbox{ind}_{D}}$;
         Set $(\mbox{ind}_{X}^{1},\mbox{ind}_{X}^{2})\leftarrow\operatorname*{arg\,max}_{(i,j)\in S\times(\bar{S}\cap\mathcal{A})}\mathcal{R}\left(\boldsymbol{x}-\boldsymbol{e}_{i}+\boldsymbol{e}_{j}\right)$ and $\boldsymbol{x}_{X}\leftarrow\boldsymbol{x}-\boldsymbol{e}_{\mbox{ind}_{X}^{1}}+\boldsymbol{e}_{\mbox{ind}_{X}^{2}}$.
         if $\max\{\mathcal{R}(\boldsymbol{x}_{D}),\mathcal{R}(\boldsymbol{x}_{X})\}\geq 1.001\times\mathcal{R}(\boldsymbol{x})$ then
              if $\mathcal{R}(\boldsymbol{x}_{D})\geq\mathcal{R}(\boldsymbol{x}_{X})$ then
                  Set $\boldsymbol{x}\leftarrow\boldsymbol{x}_{D}$ and $\mathcal{T}(\mbox{ind}_{D})\leftarrow\mathcal{T}(\mbox{ind}_{D})+1$.
              else
                  Set $\boldsymbol{x}\leftarrow\boldsymbol{x}_{X}$ and $\mathcal{T}(\mbox{ind}_{X}^{1})\leftarrow\mathcal{T}(\mbox{ind}_{X}^{1})+1$.
              end if
         else
              Break;
         end if
     end if
end while
Return: $\boldsymbol{x}$ and $\mathcal{R}(\boldsymbol{x})$.
Algorithm 3 GCN-LS Policy
Input: Indices $p_{i}$ for $i=1,2,\dots,N$.
Implement the index policy in Algorithm 1 by plugging in $p_{i}$’s. The output is denoted by $(\boldsymbol{x}_{1},\mathcal{R}_{1})$.
Implement the local-search policy in Algorithm 2 by plugging in $(\boldsymbol{x}_{1},\mathcal{R}_{1})$. The output is denoted by $(\boldsymbol{x}_{2},\mathcal{R}_{2})$
Return: $\boldsymbol{x}_{2}$ and $\mathcal{R}_{2}$.
Appendix B Discussion of Other Choice Models
In this section, we discuss a few other choice models, and provide their corresponding graph representations.
Exponomial model Alptekinoğlu and Semple (2016). In this model, the random utility of product $j$ is as follows:
$U_{j}=u_{j}-r_{j}-\xi_{j},$
where $\xi_{j}$ is an independent exponential random variables with rate $\lambda_{j}$. We also have $u_{0}=0$, $r_{0}=0$, and $U_{0}=-\xi_{0}$. Then, according to Alptekinoğlu and Semple (2016), given the assortment represented by $\boldsymbol{x}$, the choice probability of product $j$ is as follows
$\psi_{j}^{\mbox{EXP}}(\boldsymbol{x})$ $=\mathbb{P}\left(U_{j}\cdot x_{j}\geq U_{\ell}\cdot x_{\ell},\forall\ell\in[N]\cup\{0\}\right)$
$=\frac{\lambda_{j}}{L_{\delta^{-1}(j)}}\exp\left[-\sum_{i=\delta^{-1}(j)}^{N+1}\lambda_{\delta(i)}\left(u_{\delta(i)}-r_{\delta(i)}-u_{j}+r_{j}\right)\right]$
$\qquad-\sum_{l=1}^{\delta^{-1}(j)-1}\frac{\lambda_{\delta(l)}\lambda_{j}}{L_{\delta(l)}L_{\delta(l+1)}}\exp\left[-\sum_{i=l}^{N+1}\lambda_{\delta(i)}\left(u_{\delta(i)}-r_{\delta(i)}-u_{\delta(l)}+r_{\delta(l)}\right)\right],$
where $\delta$ is the ascending order of products (including the no-purchase option) by $(u_{j}-r_{j})$’s, and $L_{i}=\sum_{j=i}^{N+1}\lambda_{\delta(j)}$.
Then, we provide the graph representation under the exponomial model. In Figure 5, we use the same graph representation in Figure 1 with $K=1$. However, the feature vectors are slightly different. The feature vector of the only customer node is $[0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,1\right]$; the feature vector of constraint node $i$ is $[0,0,b_{i}]$. The feature vector of the edge from the customer node to product node $j$ is $[v_{j},\lambda_{j},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $[0,0,A_{ij}]$.
Figure 5: Graph representation of constrained assortment problem under exponomial choice model.
Nested Logit (NL) model McFadden (1980). In this model, the products are organized by nests, and customers’ selection proceeds in two stages. First, the customer chooses one of the nests or leaves without purchase. If one nest is chosen, then the customer enters the second stage and purchases one of the products in the nest or leaves without purchase. Let $v_{\ell j}$ denote the attraction value of product $j$ in nest $S_{\ell}$. In the first stage, given $L$ nests $\{S_{\ell}\}_{\ell=1}^{L}$, the choice probability of the nest $S_{\ell}$ is
$Q(S_{h}\mid\boldsymbol{x})=\frac{\left(\sum_{j\in S_{h}}x_{j}v_{hj}\right)^{\gamma_{h}}}{1+\sum_{\ell=1}^{L}\left(\sum_{j\in S_{\ell}}x_{j}v_{\ell j}\right)^{\gamma_{\ell}}},$
where $\gamma_{\ell}$ is the dissimilarity parameter of the nest $S_{\ell}$. In the second stage, given that the nest $S_{h}$ is selected, the choice probability of product $j\in S_{h}$ is
$q(j\mid S_{h},\boldsymbol{x})=\frac{x_{j}v_{hj}}{\sum_{j^{\prime}\in S_{h}}x_{j^{\prime}}v_{hj^{\prime}}}.$
Therefore, the overall choice probability of product $j$ is as follows:
$\phi_{j}^{\mbox{NL}}(\boldsymbol{x})=\sum_{\ell=1}^{L}Q(S_{\ell}\mid\boldsymbol{x})\cdot\mathbbm{1}\{j\in S_{\ell}\}\cdot q(j\mid S_{\ell},\boldsymbol{x})=\sum_{\ell=1}^{L}\frac{\left(\sum_{j\in S_{\ell}}x_{j}v_{\ell j}\right)^{\gamma_{\ell}}}{1+\sum_{\ell^{\prime}=1}^{L}\left(\sum_{j\in S_{\ell^{\prime}}}x_{j}v_{\ell^{\prime}j}\right)^{\gamma_{\ell^{\prime}}}}\cdot\frac{x_{j}v_{\ell j}}{\sum_{j^{\prime}}x_{j^{\prime}}v_{\ell j^{\prime}}}.$
As for the graph representation, we need to modify Figure 1 to include nest nodes ($\bigstar$). In Figure 6, we provide the graph representation of the problem under the NL model. Note that when connecting nest nodes and product nodes, we only connect each nest node with products in the nest. The feature vector of the only customer node is $[0,0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,0,1\right]$; the feature vector of constraint node $i$ is $[0,0,0,b_{i}]$; the feature vector of nest node $h$ is $[0,\gamma_{h},0,0]$. The feature vector of the edge from the customer node to nest node $h$ is $[1,0,0]$, the feature vector of the edge between nest node $h$ and product node $j$ is $[0,v_{hj},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $[0,0,A_{ij}]$.
Figure 6: Graph representation of constrained assortment problem under NL choice model.
Markov Chain (MC) model Blanchet et al. (2016). In this model, we assume that there are $\beta_{j}$ proportion of customers whose most preferred choice is product $j$, and $\sum_{j=0}^{N}\beta_{j}=1$. If a customer preferring product $j_{1}$ finds that the product is not available, then he switches to product $j_{2}\in[N]\cup\{0\}\setminus\{j_{1}\}$ with probability $\rho_{j_{1}j_{2}}$. If the switched product $j_{2}$ is still unavailable, then he switches to product $j_{3}\in[N]\cup\{0\}\setminus\{j_{2}\}$ with probability $\rho_{j_{2}j_{3}}$. The process continues until the customer purchases a product or switches to product 0, i.e., the no-purchase option. Note that the switching probabilities only depend on the last switched product, and hence the product switching process can be seen as a Markov chain. Let $\boldsymbol{\beta}_{S}$ denote the proportion vector of customers preferring products in $S$, and $\boldsymbol{\rho}_{(S_{1},S_{2})}$ denote the transition matrix from $S_{1}$ to $S_{2}$. Then, given the assortment $\boldsymbol{x}$, the choice probability of product $i$ is
$\phi_{j}^{\mbox{MC}}(\boldsymbol{x})=\beta_{j}+\left(\boldsymbol{\beta}_{S_{\bar{\boldsymbol{x}}}}\right)^{T}\left(\boldsymbol{I}-\boldsymbol{\rho}_{(S_{\bar{\boldsymbol{x}}},S_{\bar{\boldsymbol{x}}})}\right)^{-1}\boldsymbol{\rho}_{(S_{\bar{\boldsymbol{x}}},S_{\boldsymbol{x}}\cup\{0\})}\boldsymbol{e}_{j},$
where $S_{\boldsymbol{x}}=\{i\in[N]\mid x_{i}=1\}$, $\bar{\boldsymbol{x}}=\boldsymbol{1}-\boldsymbol{x}$, $\boldsymbol{I}$ is the identity matrix, and $\boldsymbol{e}_{j}$ is an $N$-dimension all-zero vector except that its $j$-th entry is 1.
As for the graph representation, in addition to edges in Figure 1, we introduce the product-to-product edges to represent the transition probabilities. In Figure 7, we provide the graph representation of the problem under the MC model. The feature vector of the only customer node is $[0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,1\right]$; the feature vector of constraint node $i$ is $[0,0,b_{i}]$. The feature vector of the edge from the customer node to product node $j$ is $[\beta_{j},0,0]$, the feature vector of the directed edge from product node $j_{1}$ to product node $j_{2}$ is $[0,\rho_{j_{1}j_{2}},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $[0,0,A_{ij}]$.
Figure 7: Graph representation of constrained assortment problem under MC choice model.